Anwer the questions below using the following​ information?

The growth rate of real GDP is

2.3%.

The growth rate of nominal GDP is

5.1​%.

The nominal interest rate is

7.1​%.

The real interest rate is

4.3​%.

The money supply​ (M2) is

10,612

​(in billions)

Using the information given​ above,

The growth rate of the money supply must be ​%.

​ (Round your answer to the nearesttenth.​)

The inflation rate is ​%.
Answer #1

Answer

(a)

According to Quantity theory of Money :

MV = PY

where M = Quantity of money supply, V = Velocity of Money, P = Price level and Y = Real GDP and P*Y = Nominal GDP

Formula :

% change in (A*B) = % change in A + % change in B

=> % change in (MV) = % change in (P*Y) = % change in (Nominal GDP)

=> % change in M + % change in V = % change in (Nominal GDP)

It is given that % change in (P*Y) = % change in Nominal GDP = 5.1%.

According to Quantity theory of money, Velocity of money is constant and hence % change in V = 0

=> % change in M + 0 = 5.1%

=> % change in M = 5.1%

Hence, The growth rate of the money supply must be ​5.1%.

(b)

Formula :

% change in (PY) = % change in P + % change in Y

It is given that % change in (P*Y) = % change in Nominal GDP = 5.1% and % change in Y = % change in Real GDP = 2.3%

=> 5.1% = % change in P + % change in Y + 2.3%

=> % change in P = 2.8%

As inflation rate is defined as the % increase in price level

Hence Inflation rate = % change in Price level = % change in P = 2.8%(Note P increases thus there is inflation).

Hence, The inflation rate = 2.8%.